Solve the system by elimination. If the system is inconsistent or has dependent equations, say so.
-7x + 5y = -20
-3x - 2y = 8

Answer by stanbon(75887)

1st: -7x + 5y = -20
2nd: -3x - 2y = 8
To eliminate the y-terms do as follows:
Multiply 1st thru by 2 to get:
1st: -14x+10y=-40
Multiply 2nd thru by 5 to get:
2nd: -15x-10y=40
Add these two equations to get:
-29x=0
x=0
Substitute that value into original 1st or 2nd to solve for "y".
-7(0)+5y=-20
y=-4
SOLUTION: x=0,y=4
Cheers,
stan H.